In a lottery, you pick 2 digits (from 0-9) in order, but the digits may not be repeated. A lottery ticket costs $1, and you win $40 if your ticket matches the winning numbers. What is the expected value of playing the lottery? Show your work.
Answer #1

Total number of possibilities of selecting two digits = 10*9 = 90

In this, there will be only one combo which matches the winning number. Hence,

P(Winning) = 1/90

So P(Not winning) = 1 - 1/90 = 89/90

Hence,

Expected value of playing the lottery

= $(40 - 1)*(1/90) - $1*(89/90)

= $ -0.55